good evening
Three friends A, B and C dine out and decide to split the bill according to what they order. A’s portion and B’s
portion of the bill totalled $35, while B’s portion of the bill was $5 less than C’s portion of the bill. The ratio of A’s
portion to C’s portion of the bill is 4:5. Determine the total bill amount.

Answer by Boreal(15235)

A+B=35
B+5=C
A:C=4:5
A=35-B
-
Therefore A:C=35-B:B+5 and that is 4:5
(35-B)/(B+5)=4/5
175-5B=4B+20
9B=155
B=$17.22
A=$17.78
C=$22.22
total bill amount is sum or $57.22